Two vectors of magnitude 3,4 acting at an angle 60 degree.find the resultant.

Grade:11

2 Answers

NASRUDHEEN N
6 years ago
vector A=3 vector B=4 ANGLE
NASRUDHEEN N
6 years ago
continuation of the answer
angle theta=60 degree
applying law of cosines to find the resultant magnitude R
R=sqrt(A^2+B^2+2AB cos theta)
  =sqrt(37)
R=sqrt(37)
It is not mentioned in question to find resultant direction.If it is needed apply law of sines.